# UGG-ReID: Uncertainty-Guided Graph Model for Multi-Modal Object Re-Identification

### v2

@@ 3.3.1 Uncertainty-Guided Experts Network
In the traditional MoE, the selection of experts is often based on the input content. In the application of multi-modal data, experts are usually selected independently in each modality, which makes it easy to ignore the sample noise and complementarity between modalities [27, 35]. The proposed experts network can bridge this weakness. To be specific, we first model the uncertainty of the extracted features of each modality to enhance the ability of experts to process samples with different noise levels. The multivariate Gaussian distribution that maps the features of each modality is defined as follows,
@@
However, to ensure the stability and robustness of the model expression, we do not directly use the noisy sampled feature ${z}^{m}$ in the prediction process of the final task, but use the mean as the final feature representation of the sample for the downstream decision modeling. In addition to this, to further the uncertainty modeling capabilities, we introduce the KL divergence regularization term to ensure that the sample distribution is close to the normal distribution as follows,
@@ 3.3.2 Uncertainty-Guided Routing
To better manage input features effectively, we design an uncertainty-guided routing that includes a gate mechanism to gain modal interaction by selecting different modal experts. The routing process first applies a linear transformation to the input feature. Then, using a softmax activation acts on the result of this transformation to obtain a probability score $S(\tilde{x}^{m})\in\mathbb{R}^{C}$ where $C$ denotes the number of experts. Finally, the ${TOP}_{k}$ operation selects the $k$ excellent experts from the current modality (with $k=1$) to assist the learning of other modalities, optimizing multi-modal interactions. Thus, each gate involves the total number $C+M-1$ experts, where $M$ denotes the total number of multi-modal data. Besides, to improve the experts’ ability, we further add this constraint term as follows,
@@ 3.3.3 Interactive Aggregation
The learned feature tends to be obtained by specific experts, which means that the existence of some experts can not be optimized. To solve this problem, we further add regular terms as,
@@
where $B$ denotes the batch size and ${\tilde{X}}^{m}$ is the features collection of samples in batch for the $m$-th modality. The former item refers to the proportion of samples assigned to expert $c$, and the latter item refers to the proportion of weights assigned by the router to expert $c$. $B$ denotes the batch size. Then, we aggregate interactive features via the concatenation operation as $\mathbf{\hat{z}}=[\hat{z}^{R},\hat{z}^{N},\hat{z}^{T}]$.
@@ 3.4 Train Loss
In training the proposed UGG-ReID, we combine multiple loss functions to optimize the overall framework. First, $\mathcal{L}^{m}_{c,s}$ is used to constrain the global features, which represents the sum of $\mathcal{L}^{m}_{c}$ and $\mathcal{L}^{m}_{s}$. Then, to improve the network of uncertainty-guided experts, we introduce $\mathcal{L}^{m}_{r}$, which prioritizes experts with low uncertainty by dynamically adjusting expert selection, while adopting $\mathcal{L}^{m}_{e}$ to prevent over-reliance on certain experts. Finally, cross-entropy loss $\mathcal{L}_{ce}$ and triplet loss $\mathcal{L}_{tri}$ are used to supervise the entire network. This optimization loss can be expressed by the following formula,
@@
where ${\lambda}_{1}$, ${\lambda}_{2}$ and ${\lambda}_{3}$ are the balancing coefficients this overall loss term.
@@ 4.1 Experiments Setting
Datasets. We conduct five multi-modal object ReID datasets, including two person ReID datasets (e.g., RGBNT201[9], Market1501-MM [36]) and three vehicle datasets (e.g., MSVR310 [10], RGBNT100 [8], WMVEID863 [11]). These datasets pose multi-dimensional challenges such as perspective changes and environmental disturbances, reflecting the broad applicability of this method.
@@
Methods Publication Structure RGBNT201 Market1501-MM
@@
Multi-modal HAMNet [8] AAAI20 CNN 27.7 26.3 41.5 51.7 60.0 82.8 92.5 95.0
PFNet [9] AAAI21 CNN 38.5 38.9 52.0 58.4 60.9 83.6 92.8 95.5
IEEE [36] AAAI22 CNN 46.4 47.1 58.5 64.2 64.3 83.9 93.0 95.7
TIENet [30] TNNLS25 CNN 54.5 54.4 66.3 71.1 67.4 86.1 94.1 96.0
UniCat [37] NIPSW23 ViT 57.0 55.7 - - - - - -
@@
HTT [17] AAAI24 ViT 71.1 73.4 83.1 87.3 67.2 81.5 95.8 97.8
TOP-ReID [20] AAAI24 ViT 72.2 75.2 84.9 89.4 82.0 92.4 97.6 98.6
ICPL-ReID [29] TMM25 CLIP 75.1 77.4 84.2 87.9 - - - -
PromptMA [23] TIP25 CLIP 78.4 80.9 87.0 88.9 83.6 93.3 - -
MambaPro [18] AAAI25 CLIP 78.9 83.4 89.8 91.9 84.1 92.8 97.7 98.7
DeMo [27] AAAI25 CLIP 79.7 81.8 89.4 92.5 83.6 93.1 97.5 98.7
IDEA [22] CVPR25 CLIP 80.2 82.1 90.0 93.3 - - - -
@@ 4.2 Comparison with State-of-the-Art Methods
Evaluation on Multi-modal Person ReID. We evaluate our proposed UGG-ReID on two multi-modal person ReID datasets in Table 1. We can observe that both datasets achieve state-of-the-art results for all metrics. Particularly, RGBNT201 [9] outperforms the next most popular method in the metric rank-1 by 3.4%. In addition, our method surpasses DeMo [27] in multiple metrics, highlighting the effectiveness of our architectural design. DeMo [27] relies on modality decoupling to preserve specific cues and uses attention to assign expert weights in MoE. In contrast, our proposed GPGR builds stronger modality-specific representations by incorporating epistemic uncertainty from local details. Meanwhile, the UGMoE strategy utilizes sample uncertainty to guide expert selection and applies a novel routing strategy to facilitate more effective multi-modal collaboration.
Evaluation on Multi-modal Vehicle ReID. We further conduct experiments on three multi-modal vehicle datasets, which contain challenges such as large view discrepancies and intense glare conditions, to fully validate the robustness and effectiveness of the proposed UGG-ReID. As shown in Table 2, our method maintains stable performance under various complex interference conditions, clearly demonstrating the practical effectiveness of the proposed method in enhancing the model’s generalization capability. Notably, on WMVEID863 [11] with severe dazzle interference, our method outperforms the suboptimal method by 2.8% and 3.6% in the mAP and R-1, respectively. This result further validates the robustness and effectiveness in the face of significant noise interference.
@@
Multi-modal HAMNet [8] AAAI20 CNN 27.1 42.3 74.5 93.3 45.6 48.5 63.1 68.8
PFNet [9] AAAI21 CNN 23.5 37.4 68.1 94.1 50.1 55.9 68.7 75.1
IEEE [36] AAAI22 CNN 21.0 41.0 61.3 87.8 45.9 48.6 64.3 67.9
CCNet [10] INFFUS23 CNN 36.4 55.2 77.2 96.3 50.3 52.7 69.6 75.1
@@
TOP-ReID [20] AAAI24 ViT 35.9 44.6 81.2 96.4 67.7 75.3 80.8 83.5
FACENeT [11] INFFUS25 ViT 36.2 54.1 81.5 96.9 69.8 77.0 81.0 84.2
PromptMA [23] TIP25 CLIP 55.2 64.5 85.3 97.4 - - - -
MambaPro [18] AAAI25 CLIP 47.0 56.5 83.9 94.7 69.5 76.9 80.6 83.8
DeMo [27] AAAI25 CLIP 49.2 59.8 86.2 97.6 68.8 77.2 81.5 83.8
IDEA [22] CVPR25 CLIP 47.0 62.4 87.2 96.5 - - - -
ICPL-ReID [29] TMM25 CLIP 56.9 77.7 87.0 98.6 67.2 74.0 81.3 85.6
@@ 4.3 Ablation Study
To analyze the contribution of each module, we conduct systematic ablation experiments around the two core components, UGMoE and GPGR, on RGBNT201 [9] and WMVEID863 [11]. We first use the pre-trained shared CLIP visual encoder as the baseline in Table 3, and gradually introduce each component for comparative analysis. UGMoE. Introducing the traditional Mixture of Experts (MoE) strategy to the baseline can bring some performance improvement in Table 3 (b). We further incorporate uncertainty modeling into the MoE strategy and the model performance continues to improve. Compared with the baseline, the proposed method improves mAP/R-1 by 4.8%/7.7% and 3.3%/4.6% on RGBNT201 [9] and WMVEID863 [11], respectively, which indicates that the introduction of uncertainty helps to estimate the expert’s credibility more accurately, thus realizing a more reasonable sample assignment and improving the effectiveness of multi-modal information fusion. GPGR. In the case of relying only on the above strategy for multi-modal modeling, the modal-specific information is still not fully explored. To further promote the model’s ability, we introduce GPGR to enhance the modeling ability of fine-grained local cues by using uncertainty. In the experiment, we first evaluate the effect of the standard Patch-Graph Representation (PGR) module in Table 3 (d), and then integrate the uncertainty into the network. The results show that GPGR delivers significant improvements based on the integrated UGMoE. Compared to using UGMoE alone, our method improves mAP/R-1 by 5.2%/6.4% on RGBNT201 [9] and by 2.9%/3.8% on WMVEID863 [11], respectively. This fully shows that GPGR can effectively suppress noise interference and mine richer information. In summary, our proposed method enhances features within each modality and jointly captures complementary information among multi-modal data, improving model robustness and effectiveness. More ablation is provided in the supplementary material.
@@ 4.4 Robustness Analysis
To systematically evaluate the robustness of the proposed UGG-ReID under noise interference, we inject Gaussian noise of different intensities into the RGBNT201 dataset [9], and increase the noise intensity $\varepsilon$ from 5 to 30 to generate multiple noisy versions of the dataset. Compared to the four mainstream methods (EDITOR [21], TOP-ReID [20], ICPL-ReID [29] and DeMo [27]), the proposed method maintains superior performance under all noise intensities, as shown in Fig. 3 (a). These results indicate that the method has good robustness and generalization ability.
Furthermore, we evaluate the model’s performance by injecting Gaussian noise of different intensities in the testing phase after completing training on clean data. As Fig. 3 (b), the proposed method is stable and superior to other methods under multi-level noise conditions. Meanwhile, we perform rank-list retrieval evaluations on clean data and different types of noise in Fig. 4. The results show that the proposed method achieves excellent performance under various interference conditions, which is significantly better than the advanced method DeMo [27], which fully reflects its robustness and practicability under complex perceptual interference.
@@ 4.5 Visual Results
We adopt the T-SNE method [39] to visualize the extracted features to intuitively show the feature distribution of the model under different combinations of modules. With the gradual introduction of the proposed modules, the feature distribution gradually shows obvious clustering in Fig. 5. This compact clustering structure is consistent with the performance enhancement results in Table 3, which further validates the effectiveness of our method in enhancing modal feature representation and multi-modal interaction modeling. More visualizations are provided in the supplementary material.

### v3

@@ 3.3.1 Uncertainty-Guided Experts Network
In the traditional MoE, the selection of experts is often based on the input content. In the application of multi-modal data, experts are usually selected independently in each modality, which makes it easy to ignore the sample noise and complementarity between modalities [wang2024demo, /nips/YunCPWBZXLC24]. The proposed experts network can bridge this weakness. To be specific, we first model the uncertainty of the extracted features of each modality to enhance the ability of experts to process samples with different noise levels. The multivariate Gaussian distribution that maps the features of each modality is defined as,
@@
However, to ensure the stability and robustness of the model expression, we do not directly use the noisy sampled feature ${z}^{m}$ in the prediction process of the final task, but use the mean as the final feature representation of the sample for the downstream decision modeling. In addition to this, to further the uncertainty modeling capabilities, we introduce the KL divergence [9157756] regularization term to ensure that the sample distribution is close to the normal distribution as follows,
@@ 3.3.2 Uncertainty-Guided Routing
To better manage input features effectively, we design an uncertainty-guided routing that includes a gate mechanism to gain modal interaction by selecting different modal experts. The routing process first applies a linear transformation to the input feature. Then, using a softmax activation acts on the result of this transformation to obtain a probability score $S(\tilde{x}^{m})\in\mathbb{R}^{C}$ where $C$ denotes the number of experts. Finally, the ${TOP}_{k}$ operation is employed to select the top $k$ ($k=1$) excellent experts of the current modality to learn other modalities, optimizing multi-modal interactions. Thus, each gate involves the total number C+M1C+M-1 experts, where $M$ denotes the total number of multi-modal data. Besides, to improve the experts’ ability, we further add this constraint term [10.1145/3664647.3680949] as follows,
@@ 3.3.3 Interactive Aggregation
The learned feature tends to be obtained by specific experts, which means that the existence of some experts can not be optimized. To solve this problem, we further add regular terms [10.1145/3664647.3680949] as,
@@
where $B$ denotes the batch size and ${\tilde{X}}^{m}$ is the features collection of samples in batch for the $m$-th modality. The former item refers to the proportion of samples assigned to expert $c$, and the latter item refers to the proportion of weights assigned by the router to expert $c$. Then, we aggregate interactive features via the concatenation operation as $\mathbf{\hat{z}}=[\hat{z}^{R},\hat{z}^{N},\hat{z}^{T}]$.
@@ 3.4 Train Loss
In the training of the proposed UGG-ReID, we combine multiple loss functions to optimize the overall framework. First, $\mathcal{L}^{m}_{c,s}$ is used to constrain the global features, which represents the sum of $\mathcal{L}^{m}_{c}$ and $\mathcal{L}^{m}_{s}$. Then, to improve the network of uncertainty-guided experts, we introduce $\mathcal{L}^{m}_{r}$, which prioritizes experts with low uncertainty by dynamically adjusting expert selection, while adopting $\mathcal{L}^{m}_{e}$ to prevent over-reliance on certain experts. Finally, cross-entropy loss $\mathcal{L}_{ce}$ and triplet loss $\mathcal{L}_{tri}$ are used to supervise the entire network. This optimization loss can be expressed by the following formula,
@@
where ${\lambda}_{1}$, ${\lambda}_{2}$ and ${\lambda}_{3}$ are the balancing coefficients of this overall loss term.
@@ 4.1 Experiments Setting
Datasets. We conduct five multi-modal object ReID datasets, including two person ReID datasets (e.g., RGBNT201[Zheng2021RobustMP], Market1501-MM [Wang2022InteractEA]) and three vehicle datasets (e.g., MSVR310 [ZHENG2023101901], RGBNT100 [Li2020MultiSpectralVR], WMVEID863 [ZHENG2025102800]). These datasets pose multi-dimensional challenges such as perspective changes and environmental disturbances, reflecting the broad applicability of this method.
@@
Methods Publication Structure RGBNT201 Market1501-$M$
@@
Multi-modal
HAMNet [Li2020MultiSpectralVR] AAAI20 CNN 27.7 26.3 41.5 51.7 60.0 82.8 92.5 95.0
PFNet [Zheng2021RobustMP] AAAI21 CNN 38.5 38.9 52.0 58.4 60.9 83.6 92.8 95.5
IEEE [Wang2022InteractEA] AAAI22 CNN 46.4 47.1 58.5 64.2 64.3 83.9 93.0 95.7
TIENet [Yang25TIENet] TNNLS25 CNN 54.5 54.4 66.3 71.1 67.4 86.1 94.1 96.0
UniCat [Crawford2023UniCatCA] NIPSW23 ViT 57.0 55.7 - - - - - -
@@
HTT [Wang2024HeterogeneousTT] AAAI24 ViT 71.1 73.4 83.1 87.3 67.2 81.5 95.8 97.8
TOP-ReID [Wang2023TOPReIDMO] AAAI24 ViT 72.2 75.2 84.9 89.4 82.0 92.4 97.6 98.6
ICPL-ReID [li2025icpl] TMM25 CLIP 75.1 77.4 84.2 87.9 - - - -
PromptMA [Zhang25PromptMA] TIP25 CLIP 78.4 80.9 87.0 88.9 83.6 93.3 - -
MambaPro [Wang2024MambaPro] AAAI25 CLIP 78.9 83.4 89.8 91.9 84.1 92.8 97.7 98.7
DeMo [wang2024demo] AAAI25 CLIP 79.7 81.8 89.4 92.5 83.6 93.1 97.5 98.7
IDEA [wang2025IDEA] CVPR25 CLIP 80.2 82.1 90.0 93.3 - - - -
@@ 4.2 Comparison with State-of-the-Art Methods
Evaluation on Multi-modal Person ReID. We evaluate our proposed UGG-ReID on two multi-modal person ReID datasets in Table 1. We can observe that both datasets achieve state-of-the-art results for all metrics. Particularly, RGBNT201 [Zheng2021RobustMP] outperforms the next most popular method in the metric rank-1 by 3.4%. In addition, our method surpasses DeMo [wang2024demo] in multiple metrics, highlighting the effectiveness of our architectural design. DeMo [wang2024demo] relies on modality decoupling to preserve specific cues and uses attention to assign expert weights in MoE. In contrast, our proposed GPGR builds stronger modality-specific representations by incorporating aleatoric uncertainty from local details. Meanwhile, the UGMoE strategy utilizes sample uncertainty to guide expert selection and applies a novel routing strategy to facilitate more effective multi-modal collaboration.
Evaluation on Multi-modal Vehicle ReID. We further conduct experiments on three multi-modal vehicle datasets, which contain challenges such as large view discrepancies and intense glare conditions, to fully validate the robustness and effectiveness of the proposed UGG-ReID. As shown in Table 2, our method maintains stable performance under various complex interference conditions, clearly demonstrating the practical effectiveness of the proposed method in enhancing the model’s generalization capability. Notably, on WMVEID863 [ZHENG2025102800] with severe dazzle interference, our method outperforms the suboptimal method by 2.8% and 3.6% in the mAP and R-1, respectively. This result further validates the robustness and effectiveness in the face of significant noise interference.
@@
Multi-modal
HAMNet [Li2020MultiSpectralVR] AAAI20 CNN 27.1 42.3 74.5 93.3 45.6 48.5 63.1 68.8
PFNet [Zheng2021RobustMP] AAAI21 CNN 23.5 37.4 68.1 94.1 50.1 55.9 68.7 75.1
IEEE [Wang2022InteractEA] AAAI22 CNN 21.0 41.0 61.3 87.8 45.9 48.6 64.3 67.9
CCNet [ZHENG2023101901] INFFUS23 CNN 36.4 55.2 77.2 96.3 50.3 52.7 69.6 75.1
@@
TOP-ReID [Wang2023TOPReIDMO] AAAI24 ViT 35.9 44.6 81.2 96.4 67.7 75.3 80.8 83.5
FACENeT [ZHENG2025102800] INFFUS25 ViT 36.2 54.1 81.5 96.9 69.8 77.0 81.0 84.2
PromptMA [Zhang25PromptMA] TIP25 CLIP 55.2 64.5 85.3 97.4 - - - -
MambaPro [Wang2024MambaPro] AAAI25 CLIP 47.0 56.5 83.9 94.7 69.5 76.9 80.6 83.8
DeMo [wang2024demo] AAAI25 CLIP 49.2 59.8 86.2 97.6 68.8 77.2 81.5 83.8
IDEA [wang2025IDEA] CVPR25 CLIP 47.0 62.4 87.2 96.5 - - - -
ICPL-ReID [li2025icpl] TMM25 CLIP 56.9 77.7 87.0 98.6 67.2 74.0 81.3 85.6
@@ 4.3 Ablation Study
To analyze the contribution of each module, we conduct systematic ablation experiments around the two core components, UGMoE and GPGR, on RGBNT201 [Zheng2021RobustMP] and WMVEID863 [ZHENG2025102800]. We first use the pre-trained shared CLIP visual encoder as the baseline in Table 3, and gradually introduce each component for comparative analysis. UGMoE. Introducing the traditional Mixture of Experts (MoE) strategy to the baseline can bring some performance improvement in Table 3 (b). We further incorporate uncertainty modeling into the MoE strategy and the model performance continues to improve. Compared with the baseline, the proposed method improves mAP/R-1 by 4.8%/7.7% and 3.3%/4.6% on RGBNT201 [Zheng2021RobustMP] and WMVEID863 [ZHENG2025102800], respectively, which indicates that the introduction of uncertainty helps to estimate the expert’s credibility more accurately, thus realizing a more reasonable sample assignment and improving the effectiveness of multi-modal information fusion. GPGR. In the case of relying only on the above strategy for multi-modal modeling, the modal-specific information is still not fully explored. To further promote the model’s ability, we introduce GPGR to enhance the modeling ability of fine-grained local cues by using uncertainty. In the experiment, we first evaluate the effect of the standard Patch-Graph Representation (PGR) module in Table 3 (d), and then integrate the uncertainty into the network. The results show that GPGR delivers significant improvements based on the integrated UGMoE. Compared to using UGMoE alone, our method improves mAP/R-1 by 5.2%/6.4% on RGBNT201 [Zheng2021RobustMP] and by 2.9%/3.8% on WMVEID863 [ZHENG2025102800], respectively. This fully shows that GPGR can effectively suppress noise interference and mine richer information. In summary, our proposed method enhances features within each modality and jointly captures complementary information among multi-modal data, improving model robustness and effectiveness. More ablation is provided in the supplementary material.
@@ 4.4 Robustness Analysis
To systematically evaluate the robustness of the proposed UGG-ReID under noise interference, we inject Gaussian noise of different intensities into the RGBNT201 dataset [Zheng2021RobustMP], and increase the noise intensity $\varepsilon$ from 5 to 30 to generate multiple noisy versions of the dataset. Compared to the four mainstream methods (EDITOR [10654953], TOP-ReID [Wang2023TOPReIDMO], ICPL-ReID [li2025icpl] and DeMo [wang2024demo]), the proposed method maintains superior performance under all noise intensities, as shown in Fig. 3 (a). These results indicate that the method has good robustness and generalization ability.
Furthermore, we evaluate the model’s performance by injecting Gaussian noise of different intensities in the testing phase after completing training on clean data. As Fig. 3 (b) shows, the proposed method is stable and superior to other methods under multi-level noise conditions. Meanwhile, we perform rank-list retrieval evaluations on clean data and different types of noise in Fig. 4. The results show that the proposed method achieves excellent performance under various interference conditions, which is significantly better than the advanced method DeMo [wang2024demo], which fully reflects its robustness and practicability under complex perceptual interference.
@@
Table 4: Expert-balanced analysis of each modality on WMVEID863.
M Score $E_{1}$ $E_{2}$ $E_{3}$ $E_{4}$
R Uncer. 0.31 0.27 0.30 0.12
Gate 0.07 0.12 0.07 0.74
N Uncer. 0.24 0.24 0.27 0.24
Gate 0.28 0.26 0.18 0.28
T Uncer. 0.28 0.26 0.29 0.18
Gate 0.17 0.21 0.14 0.48
Table 5: Efficiency analysis and accuracy comparison with SOTA methods on RGBNT201.
RGBNT201 Params(M) FLOPs(G) FPS mAP R-1
TOP-ReID [Wang2023TOPReIDMO] 324.5 35.5 398.9 72.2 75.2
EDITOR [10654953] 119.3 40.8 335.1 66.7 68.7
PromptMA [Zhang25PromptMA] 107.9 36.2 343.5 78.4 80.9
MambaPro [Wang2024MambaPro] 74.8 52.4 243.2 78.9 83.4
DeMo [wang2024demo] 98.8 35.1 403.6 79.7 81.8
IDEA [wang2025IDEA] 91.7 43.7 299.5 80.2 82.1
UGG-ReID 103.2 35.0 371.4 81.2 86.8
4.5 Expert Balanced Analysis
We show the router average gate scores and uncertainty scores of all test samples for each modality on WMVEID863 in Table 5. We observe that N and T modalities show a more balanced expert activation distribution, while Expert 4 in the R modality has a heavy weight when the uncertainty is low. These results indicate that the routing mechanism dynamically adjusts expert allocation according to different modalities and sample uncertainty.
4.6 Efficiency Analysis
To further validate the effectiveness of our proposed UGG-ReID, we conduct an efficiency analysis, as shown in Table 5. We evaluate the inference speed of each method on the RGBNT201 dataset, using Frames Per Second (FPS) as the evaluation metric. As shown in the Table 5, UGG-ReID reaches an inference speed of 371.4 FPS while maintaining a relatively low parameter count of 103.2 million and a computational cost of 35.0G FLOPs. This speed is only slightly lower than that of the lighter DeMo [wang2024demo], and is significantly higher than most mainstream approaches, including MambaPro [Wang2024MambaPro] and EDITOR [10654953]. Notably, despite its high efficiency, UGG-ReID still achieves excellent results.
Figure 5: Visualization results of the (a) Input image. (b) Baseline. (c) UGG-ReID (Ours).
@@ 4.7 Visual Results
Multi-modal Activation Maps: We adopt the Grad-CAM method [8237336] to visualize the extracted features of each modality. Samples with backpacks are above the dotted line, and samples without backpacks are below the dotted line. The proposed UGG-ReID consistently attends to more semantically rich object regions compared to the baseline model that lacks our module, as shown in Fig. 5. Notably, under challenging conditions such as motion blur or occlusion, our method still effectively focuses on key target features, further validating its robustness and discriminative power. This suggests that our method effectively captures the diversity of multi-modal information through uncertainty-guided multi-modal local and sample-level joint learning, which significantly improves the robustness and the performance of object re-identification.
Multi-modal Feature Distributions: We adopt the T-SNE method [Kobak2021InitializationIC] to visualize the extracted features to intuitively show the feature distribution of the model under different combinations of modules. With the gradual introduction of the proposed modules, the feature distribution gradually shows obvious clustering in Fig. 6. This compact clustering structure is consistent with the performance enhancement results in Table 3, which further validates the effectiveness of our method in enhancing modal feature representation and multi-modal interaction modeling. More visualizations are provided in the supplementary material.